# Pricing AI Model Accuracy

### v1

@@
Abstract: This paper examines the market for AI models in which firms compete to provide accurate model predictions and consumers exhibit heterogeneous preferences for model accuracy. We develop a consumer-firm duopoly model to analyze how competition affects firms’ incentives to improve model accuracy. Each firm aims to minimize its model’s error, but this choice can often be suboptimal. Counterintuitively, we find that in a competitive market, firms that improve overall accuracy do not necessarily improve their profits. Rather, each firm’s optimal decision is to invest further on the error dimension where it has a competitive advantage. By decomposing model errors into false positive and false negative rates, firms can reduce errors in each dimension through investments. Firms are strictly better off investing on their superior dimension and strictly worse off with investments on their inferior dimension. Profitable investments adversely affect consumers but increase overall welfare.
@@ 1 Introduction
In the artificial intelligence (AI) and machine learning (ML) industry, an AI model supplier’s central goal is to minimize its model’s error. This error metric can be very subjective to the task at hand; generative AI models may attempt to produce output strongly resembling the user’s intended requests, whereas traditional classification models may try to minimize incorrect predictions on test data. In all such cases, though, firms typically try to minimize some corresponding measure of overall error as a proxy for model performance.
@@
While this goal is theoretically optimal, achieving a minimal error rate can be both costly and impractical for firms (Zhang et al. (2019)). Plus, improving multiple error dimensions simultaneously can be expensive and inefficient. In addition, an improvement in one dimension may lead to a worsening in another dimension, inducing an inverse trade-off across dimensions (Yaniv and Foster (1995)). A typical example of this is binary classification, where directly minimizing both Type I and Type II errors with a fixed sample size is typically infeasible (Mudge et al. (2012)). Assuming a fixed total level of error, designing a model with fewer false positive (FP) errors would shrink the critical region but also simultaneously lead to an increase in false negative (FN) error risk, and vice versa. As a result, model designers have to prioritize between FP and FN errors based on their expectations about consumer preferences. A standard statistical approach is fixing one dimension and optimizing the other (Maier and Lakens (2022)).
@@
On the demand side, consumers often care less about the total model error and more about the specific types of errors a model makes, as some errors may be more costly to them than others. The bias-variance trade-off is an example of this inverse relation in model performance; lowering bias by improving model complexity typically leads to higher variance (Geman et al. (1992)). One specific measure of model performance that consumers often consider is based on FP and FN rates (Cappelen et al. (2023)). Different consumers exhibit varying levels of sensitivities to FP and FN rates, depending on their specific use cases and their associated cost sensitivities to FP and FN errors.
@@
A common pitfall of generative AI models is hallucination, where the model produces entirely incorrect outputs. Hallucination rates, or more broadly FP rates, of the most popular models like ChatGPT remain quite significant (Rawte et al. (2023)). Users employing AI for creative writing may have a completely different tolerance for hallucination compared to someone using the same model for disease diagnosis. This consumer heterogeneity can be captured by different sensitivities toward hallucination rates. A scientific user may prefer a non-informative model output (i.e., “I don’t know”) as opposed to a model hallucination.
@@
This implicit trade-off between FP and FN rates perceived by consumers, along with the difficulties associated with pure error reduction, leads to product differentiation by firms. Firms specialize through investments towards reducing error rates on certain dimensions to maximize profits. As a result, firms try to adjust their FP and FN rates in a way that increases profits, attracting more consumers while also maintaining a sufficiently high price. In sum, two relevant phenomena are present in the AI model market: (i) heterogeneous cost sensitivity levels among consumers and (ii) an implicit firm-level trade-off between FP and FN rates.
@@
In this paper, we develop an economic model that captures both these phenomena and produces a somewhat counter-intuitive result: firms should invest in growing their competitive advantage rather than closing the gap on their inferior error dimension. This result stems from the economic intuition that closing the gap will have a minimal positive effect on market share but a greater negative effect on equilibrium prices, resulting in the firm being worse off. We develop a traditional firm-consumer market model where firms are endowed with an initially fixed error rate, and they can then make investments toward improving either their FP or FN rate. Consumers exhibit heterogeneity through negatively correlated cost sensitivities towards FP and FN rates, which generates a natural trade-off between the two error types.
@@
We define the static version of the model in Section 3, where firms are endowed with fixed FP and FN rates, and equilibrium prices are determined endogenously. Within this fixed error rate framework, we introduce heterogeneity in the cost sensitivities of consumers. In particular, we consider perfectly positively correlated cost sensitivities, which simplifies to the Hotelling model. We then consider the more relevant case of perfectly negatively correlated sensitivities and analyze the resulting equilibria. Finally, we consider a mixture of positively and negatively correlated consumer sensitivities in the market.
Naturally, total firm revenue is maximized when all consumers have negatively correlated sensitivities, as firms are able to differentiate effectively and charge higher prices to attract consumers that match their cost sensitivities. We observe that the general model results hold under this additional heterogeneity as long as there is a non-trivial proportion of the market with negatively correlated cost sensitivities.
@@
Model dynamics are introduced in Section 4, where we consider the stability of the equilibria presented in Section 2. We show that the initial equilibrium points are not stable; profitable deviations often exist for both firms. The equilibria can be categorized as cases of split domination, where each firm dominates the other on one dimension, versus strict domination, where one firm dominates the other on both dimensions. Under split domination, we present our main result: firms should invest on the error dimension in which they are strictly superior. We also establish that such profitable investments for firms adversely affect overall consumer welfare, as increased prices result in a higher average cost per consumer.
@@ 2 Related Literature
Our work builds upon past research at the intersection of machine learning and strategic behavior, focusing on how competition affects efficiency and the resulting equilibria. Previous literature has extensively analyzed competition in such settings. For example, Immorlica et al. (2011) propose an optimization-based framework for strategic competition, but does not include prices. Feng et al. (2022) examine the bias-variance trade-off in competitive environments, demonstrating a tendency for firms to favor variance-based errors over bias-based errors. Similarly, Jagadeesan et al. (2024) propose a classification model under which competition can potentially lead to worse overall model accuracy. Other studies adopt more information-driven approaches to competition (Liang (2020)) and explore coordination among firms to learn about consumer types (Gradwohl and Tennenholtz (2023)).
@@
This paper differs from the related literature along multiple dimensions. First, we focus explicitly on competition over false positive (FP) and false negative (FN) rates rather than on more general model properties like the bias-variance trade-off. While the bias-variance framework is instrumental for understanding model complexity and generalization (Feng et al. (2022)), the FP/FN trade-off is more specific to classification tasks. Jagadeesan et al. (2024) examines competition based on accuracy measured with Bayes risk, but firms do not choose prices. Classification problems are central to many real-world applications, such as medical diagnostics, fraud detection, and content filtering, where FP and FN errors carry distinct and significant costs. By concentrating on FP and FN rates, we enable a more granular analysis of error distributions and consumer preferences.
@@
Unlike Feng et al. (2022), where firms optimize directly over bias- and variance-based errors, the FP and FN rates in our model simultaneously determine both the first and second moments of the cost function. Consumers care about the expected cost of a mistake, and a lower expected cost does not necessarily imply a higher variance. This approach is particularly effective for modeling consumer behavior, as it leverages the discrete error distributions unique to classification tasks, simplifying the analysis of how different error types impact user satisfaction.
@@
Second, we extend the existing literature by incorporating consumer market share and pricing dynamics into the model. A significant portion of prior research considers competition exclusively between firms influenced by exogenous information or static states without explicitly modeling market mechanisms. By introducing a consumer distribution and corresponding market share, we allow firms to compete not only by optimizing their error rates but also by strategically setting prices to attract consumers. In our model, consumers choose between firms based on their individual sensitivities to FP and FN errors as well as the prices charged by each firm. This resulting structure relies on market clearing conditions and endogenous price equilibria. Firms must balance the trade-off between improving model accuracy (reducing FP and FN rates) and setting competitive prices to capture a larger market share. By explicitly modeling these interactions, we link firms’ algorithmic decisions under competition to pricing and capturing market share.
@@
Third, we introduce heterogeneity among both consumers and firms. Empirical evidence suggests that consumers exhibit varying sensitivities to FP and FN errors (Cappelen et al. (2023)). We formally model this heterogeneity using FP and FN cost parameters, accounting for cardinal differences, the magnitude of costs associated with FP and FN errors for each consumer, and ordinal differences, the relative importance or priority that consumers assign to FP versus FN errors. Similarly, we incorporate firm heterogeneity through varying error endowments. Based on their initial endowments, firms may specialize in optimizing for different error types or target specific consumer segments, leading to potentially diverse equilibria.
@@ 3 The Model
Firms are endowed with a false positive (FP) rate and a false negative (FN) rate. We assume that $FP,FN\geq 0$ for all firms and assume that product differentiation exists, so no two firms have the same pair of FP and FN rates. A firm’s FP and FN rates are public information to consumers and other firms. We also assume that firms’ FN rates are less volatile than FP rates, meaning differences in FN rates across firms are smaller than differences in FP rates: $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$. This assumption imposes an ordering on the volatility of FP and FN rates; however, the model and presented results all hold under the reverse ordering as well.
Each firm $j$ chooses a non-negative price $p_{j}\in[0,\bar{p}]$ to maximize revenues, and market share to each firm is calculated as a proportion of the unit interval $[0,1]$. We assume firms do not incur any direct costs from producing their model. For simplicity, we assume a duopoly model structure with two competing firms in the market.
@@
[…]
@@ 3.1 Perfectly Positively Correlated Cost Sensitivities
We now impose structure on consumers’ cost sensitivities, first considering the case where $\alpha=\beta$, implying perfectly positively correlated consumer sensitivities to FP and FN rates. Under this structure, the model simplifies to the Hotelling model. For conciseness, we define $\mathcal{F}_{j}$ as the total error rates for firm $j$: $\mathcal{F}_{j}=FP_{j}+FN_{j}$. The consumer’s cost function $c_{i}(j)$ then simplifies to:
@@
The resulting equilibrium prices depend solely on each firm’s total error rates $\mathcal{F}_{j}$. In particular, the firm with a lower total error chooses a price twice as high as the other firm. When both firms have the same total error rates, they charge the same price and choose the highest possible price they can charge: $\bar{p}$. Precise derivations are provided in A.1.
@@
If $\textit{If }\mathcal{F}_{1}-\mathcal{F}_{2}>0:\hfil\hskip 56.9055pt&\textit{If }\mathcal{F}_{1}-\mathcal{F}_{2}<0:\hfil\hskip 56.9055pt&\textit{If }\mathcal{F}_{1}-\mathcal{F}_{2}=0:\\ p_{1}=\frac{1}{3}(\mathcal{F}_{1}-\mathcal{F}_{2})\hfil\hskip 56.9055pt&p_{1}=\frac{2}{3}(\mathcal{F}_{2}-\mathcal{F}_{1})\hfil\hskip 56.9055pt&p_{1}=\bar{p}\\ p_{2}=\frac{2}{3}(\mathcal{F}_{1}-\mathcal{F}_{2})\hfil\hskip 56.9055pt&p_{2}=\frac{1}{3}(\mathcal{F}_{2}-\mathcal{F}_{1})\hfil\hskip 56.9055pt&p_{2}=\bar{p}\end{array}$
@@ 3.2 Perfectly Negatively Correlated Cost Sensitivities
To proceed, we specify an ordering on the FP and FN rates of both firms. We consider two cases: strict domination and split domination. Under split domination, each firm dominates the other on one dimension, leading to non-zero prices for both firms. Under strict domination, one firm has strictly lower FP and FN rates than the other and thus effectively dominates the other firm. We show that the inferior firm always exits the market, and the dominating firm operates as a monopoly. Note that ordering on FP and FN rates does not affect the model in 2.1, as no distinction exists between FP and FN rates, so only total error is relevant.
@@ 3.2.1 Split Domination
Figure 1: Price equilibria under split domination with varying fixed FP and FN rate values
@@ 3.2.2 Strict Domination
Using the assumption that volatility in FN rates is smaller than the volatility in FP rates: $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$, firm 2 exits the market, and firm 1 operates as a monopoly.
@@
In particular, $p_{1}^{*}=\frac{FN_{2}-FN_{1}}{2}$ and $p_{2}^{*}=0$. As the difference in FN rates between firm 1 and firm 2 becomes smaller, firm 1 charges a lower and lower price in equilibrium to keep firm 2 out of the market. Thus, the equilibrium price $p_{1}$ is increasing in the FP and FN rate differences between firms. Choosing a higher $p_{1}$ would give firm 2 a non-zero market share, which would lower firm 1’s revenue. The reaction functions are identical to the split domination case above, but the lower bound of 0 becomes more prevalent. More details are provided in A.3, and we present a couple of examples in Figure 2.
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111(0.1, 0.1)$p_{1}$$p_{2}$$p_{1}=\max(0,0.025+0.5p_{2})$$p_{2}=\max(0,-0.05+0.5p_{1})$
(a) $FP_{1}=0.1,FP_{2}=0.2,FN_{1}=0.05,FN_{2}=0.1$; Monopoly with small $p_{1}^{*}=0.025$
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111(0.15, 0)$p_{1}$$p_{2}$$p_{1}=\max(0,0.15+0.5p_{2})$$p_{2}=\max(0,-0.2+0.5p_{1})$
(b) $FP_{1}=0.1,FN_{1}=0.05,FP_{2}=0.5,FN_{2}=0.35$; Monopoly with larger $p_{1}^{*}=0.15$
Figure 2: Price equilibria under strict domination with varying fixed FP and FN rate values

### v2

@@
Abstract: This paper examines the market for AI models in which firms compete to provide accurate model predictions and consumers exhibit heterogeneous preferences for model accuracy. We develop a consumer-firm duopoly model to analyze how competition affects firms’ incentives to improve model accuracy. Each firm aims to minimize its model’s error, but this choice can often be suboptimal. Counterintuitively, we find that in a competitive market, firms that improve overall accuracy do not necessarily improve their profits. Rather, each firm’s optimal decision is to invest further on the error dimension where it has a competitive advantage. By decomposing model errors into false positive and false negative rates, firms can reduce errors in each dimension through investments, and they are strictly better off investing on their superior dimension and strictly worse off with investments on their inferior dimension. Profitable investments adversely affect consumers but increase overall market welfare.
@@ 1 Introduction
In the artificial intelligence (AI) and machine learning (ML) industry, an AI model supplier’s central goal is to minimize its model’s error. This error metric can be very task-dependent; generative AI models prioritize producing outputs strongly resembling the user’s intended requests, whereas traditional classification models may try to minimize incorrect predictions on test data. In all such cases, though, firms typically try to minimize some corresponding measure of overall error as a proxy for model performance.
@@
While this goal is theoretically optimal, achieving a minimal error rate can be both costly and impractical for firms (Zhang et al., 2019). Improving multiple error dimensions simultaneously can be expensive and inefficient. In addition, an improvement in one dimension may lead to a worsening in another dimension, inducing an inverse trade-off across dimensions (Yaniv and Foster, 1995). A typical example of this is binary classification, where directly minimizing both Type I and Type II errors with a fixed sample size is typically infeasible (Mudge et al., 2012). Assuming a fixed total level of error, designing a model with fewer false positive (FP) errors would shrink the rejection region but also simultaneously lead to an increase in false negative (FN) error risk, and vice versa. As a result, model designers have to prioritize between FP and FN errors based on their expectations about consumer preferences. A standard statistical approach is fixing one dimension and optimizing the other (Maier and Lakens, 2022).
@@
On the demand side, consumers often care less about the total model error and more about the specific types of errors a model makes, as some errors may be more costly to them than others. The bias-variance trade-off is an example of this inverse relation in model performance; lowering bias by improving model complexity typically leads to higher variance (Geman et al., 1992). One specific measure of model performance that consumers often consider is based on FP and FN rates (Cappelen et al., 2023). Different consumers exhibit varying levels of sensitivities to FP and FN rates, depending on their specific use cases and their associated cost sensitivities to FP and FN errors.
@@
A prevalent shortcoming with generative AI models is hallucination, where the model generates entirely incorrect outputs. Hallucination rates, or more broadly FP rates, of the most popular models like ChatGPT remain quite significant (Rawte et al., 2023). Users employing AI for creative writing may have a completely different tolerance for hallucination compared to someone using the same model for disease diagnosis. This consumer heterogeneity can be captured by different sensitivities toward hallucination rates. For example, a scientific user may prefer a non-informative model output (i.e., “I don’t know”) as opposed to a model hallucination.
As a motivating example, consider two model users, Company A and Company B, both searching for new employees but struggling to choose new hires. Each company employs an AI-based labor screening model that, for every candidate, outputs either 1 (recommended hire) or 0 (recommended reject). The model therefore has two potential kinds of errors: false positives —recommending a hire who is actually unsuitable, and false negatives —rejecting a candidate who would have been a strong hire. Suppose that Company A is a large tech firm, and thus can tolerate a higher volume of applicants; its main concern is missing strong candidates, so false negatives are especially costly. Company B, on the other hand, is a small research lab with very few positions; its main concern is the disruption caused by a bad hire, so false positives are more costly.
@@
This implicit trade-off between FP and FN rates perceived by consumers, along with the difficulties associated with pure error reduction, leads to product differentiation by model suppliers. Such firms specialize through investments towards reducing error rates on certain dimensions to maximize profits; with limited resources, they must choose whether to spend on reducing their FP/FN error rates or expanding their training dataset which reduces the marginal cost of future error improvements. As a result, these supplying firms try to invest in a way that increases profits, attracting more consumers while also maintaining a sufficiently high price. In sum, two relevant phenomena are present in the AI model market: (i) heterogeneous cost sensitivity levels among consumers and (ii) an implicit firm-level trade-off between FP and FN rates.
@@
In this paper, we develop an economic model that captures both these phenomena and produces a somewhat counter-intuitive result: firms should invest in growing their competitive advantage rather than closing the gap on their inferior error dimension. This result stems from the economic intuition that closing the gap will have a minimal positive effect on market share but a greater negative effect on equilibrium prices, resulting in the firm being worse off. We develop a firm-consumer market model where firms are endowed with an initially fixed error rate, and they can then make investments toward improving either their FP or FN rate. Consumers exhibit heterogeneity through negatively correlated cost sensitivities towards FP and FN rates, which generates a natural trade-off between the two error types.
@@
We define the static version of the model in Section 3, where firms are endowed with fixed FP and FN rates, and equilibrium prices are determined endogenously. Within this fixed error rate framework, we introduce heterogeneity in the cost sensitivities of consumers. Under perfectly positively correlated cost sensitivities, our model simplifies to an extension of the standard Hotelling model. We then consider perfectly negatively correlated sensitivities and analyze the resulting equilibria in the market. Naturally, total firm revenue is maximized when all consumers have negatively correlated sensitivities, as firms are able to differentiate effectively and charge higher prices to attract consumers that match their cost sensitivities.
@@
Model dynamics are introduced in Section 4, where we consider the stability of the equilibria presented in Section 3. Firms can invest in two ways: directly improving their FP and FN error rates or increasing their firm size, and implicitly the size of their training dataset. We show that the initial market clearing prices are not equilibria; profitable deviations often exist for both firms. The equilibria can be categorized into cases of split domination, where each firm dominates the other on one dimension, versus strict domination, where one firm dominates the other on both dimensions. Under split domination, we present our main result: firms should invest on the error dimension in which they are strictly superior.
We first consider the case where firms do not face direct costs and show that even in such a simplified case, firms should never invest to close the gap on their inferior dimension. We then introduce an AI-inspired cost function and show that the optimal investment decisions can be explicitly tied to the firm’s scalability and growth potential. We also establish that such profitable investments for firms adversely affect overall consumer welfare, as increased prices result in a higher average cost per consumer.
@@ 2 Related Literature
Our work builds upon past research at the intersection of machine learning and strategic behavior, focusing on how competition affects efficiency and the resulting equilibria. Previous literature has extensively analyzed competition in such settings. For example, Immorlica et al. (2011) propose an optimization-based framework for strategic competition, but does not include prices. Feng et al. (2022) examine the bias-variance trade-off in competitive environments, demonstrating a tendency for firms to favor variance-based errors over bias-based errors. Similarly, Jagadeesan et al. (2024) propose a classification model under which competition can potentially lead to worse overall model accuracy. Other studies adopt more information-driven approaches to competition (Liang, 2020) and explore coordination among firms to learn about consumer types (Gradwohl and Tennenholtz, 2023).
@@
This paper differs from the related literature along multiple dimensions. First, we focus explicitly on competition over false positive (FP) and false negative (FN) rates rather than on the bias-variance trade-off. While the bias-variance framework is instrumental for understanding model complexity and generalization (Feng et al., 2022), the FP/FN trade-off is more specific to classification tasks. Jagadeesan et al. (2024) examines competition based on accuracy measured with Bayes risk, but firms do not choose prices. Classification problems are central to many real-world applications, such as medical diagnostics, fraud detection, and content filtering, where FP and FN errors carry distinct and significant costs. By concentrating on FP and FN rates, we enable a more granular analysis of error distributions and consumer preferences.
@@
Unlike Feng et al. (2022), where firms optimize directly over bias and variance-based errors, the FP and FN rates in our model simultaneously determine both the first and second moments of the cost function. Consumers care about the expected cost of a mistake, and a lower expected cost does not necessarily imply a higher variance. This approach is particularly effective for modeling consumer behavior, making it easier to see how different types of mistakes affect consumer satisfaction.
@@
Second, we extend the existing literature by incorporating consumer market share and pricing dynamics into the model. By introducing a consumer distribution and corresponding market share, we allow firms to compete not only by optimizing their error rates but also by strategically setting prices to attract consumers. In our model, consumers choose between firms based on their individual sensitivities to FP and FN errors as well as the prices charged by each firm. This resulting structure relies on market clearing conditions and endogenous price equilibria. Firms must balance the trade-off between improving model accuracy (reducing FP and FN rates) and setting competitive prices to capture a larger market share. By explicitly modeling these interactions, we link firms’ algorithmic decisions under competition to pricing and capturing market share.
@@
Third, we introduce heterogeneity among both consumers and firms. Empirical evidence suggests that consumers exhibit varying sensitivities to FP and FN errors (Cappelen et al., 2023). We formally model this heterogeneity using FP and FN cost parameters, accounting for cardinal differences, the magnitude of costs associated with FP and FN errors for each consumer, and ordinal differences, the relative importance or priority that consumers assign to FP versus FN errors. Similarly, we incorporate firm heterogeneity through varying error endowments. Based on their initial endowments, firms may specialize in optimizing for different error types or target specific consumer segments, leading to potentially diverse equilibria.
@@ 3 The Model
Firms are endowed with a false positive (FP) rate and a false negative (FN) rate. We assume that $FP,FN\geq 0$ for all firms and assume that there is product differentiation, so that no two firms have the same pair of FP and FN rates. A firm’s FP and FN rates are public information to consumers and other firms. We also assume that firms’ FN rates are less volatile than FP rates, meaning differences in FN rates across firms are smaller than differences in FP rates: $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$. This assumption imposes an ordering on the volatility of FP and FN rates; however, the model and presented results all hold under the reverse ordering as well.
Each firm $j$ chooses a non-negative price $p_{j}\in[0,\bar{p}]$ to maximize revenues, and market share to each firm is calculated as a proportion of the unit interval $[0,1]$. For simplicity, we assume a duopoly model structure with two competing firms in the market. We assume firms do not incur any direct costs from purely producing their model; however, they face direct investment and development costs dependent on their FP and FN rates as well as a underlying firm size parameter. We discuss this more in detail in Section 4.
@@
$\underset{j}{\text{min}}$ $c_{i}(j)\hskip 5.69046pt\Rightarrow$ $\underset{j}{\text{min}}$ $(\alpha_{i}\times FP_{j})+(\beta_{i}\times FN_{j})+p_{j}.$
@@ 3.1 Perfectly Positively Correlated Cost Sensitivities
We now impose structure on consumers’ cost sensitivities, first considering the case where $\alpha=\beta$, implying perfectly positively correlated consumer sensitivities to FP and FN rates. Under this structure, the model simplifies to the Hotelling model (Hotelling, 1981; d’Aspremont et al., 1979). For conciseness, we define $\mathcal{F}_{j}$ as the total error rates for firm $j$: $\mathcal{F}_{j}=FP_{j}+FN_{j}$. The consumer’s cost function $c_{i}(j)$ then simplifies to:
@@
[…]
The resulting equilibrium prices depend solely on each firm’s total error rates $\mathcal{F}_{j}$. The firm with lower overall error rates charges a higher price in the market equilibrium as its better product appeals to a larger market share. In particular, the firm with a lower total error chooses a price twice as high as the other firm. When both firms have the same total error rates, they charge the same price and choose the highest possible price they can charge: $\bar{p}$. Precise derivations are provided in Appendix A.1.
@@ 3.2 Perfectly Negatively Correlated Cost Sensitivities
To proceed, we specify an ordering on the FP and FN rates of both firms. We consider two cases: strict domination and split domination. Under split domination, each firm dominates the other on one dimension, leading to non-zero prices for both firms. Under strict domination, one firm has strictly lower FP and FN rates than the other and thus effectively dominates the other firm. We show that the inferior firm always exits the market, and the dominating firm operates as a monopoly. Note that ordering on FP and FN rates does not affect the model in Section 3.1, as no distinction exists between FP and FN rates and, therefore, only total error is relevant.
@@ 3.2.1 Split Domination
Figure 1: Price equilibria under split domination with varying fixed FP and FN rate values.
@@ 3.2.2 Strict Domination
Using the assumption that FN rates are less volatile than FP rates: i.e. $|FN_{2}-FN_{1}|<|FP_{2}-FP_{1}|$, firm 2 exits the market, and firm 1 operates as a monopoly.
@@
In particular, $p_{1}^{*}=\frac{FN_{2}-FN_{1}}{2}$ and $p_{2}^{*}=0$. As the difference in FN rates between firm 1 and firm 2 becomes smaller, firm 1 charges a lower and lower price in equilibrium to keep firm 2 out of the market. Thus, the equilibrium price $p_{1}$ is increasing in the FP and FN rate differences between firms. Choosing a higher $p_{1}$ would give firm 2 a non-zero market share, which would lower firm 1’s revenue. The reaction functions are identical to the split domination case above, but the lower bound of 0 becomes more prevalent. More details are provided in Appendix A.3, and we present a couple of examples in Figure 2.
[…]
(a) $FP_{1}=0.1,FP_{2}=0.2,FN_{1}=0.05,FN_{2}=0.1$; Monopoly with small $p_{1}^{*}=0.025$.
[…]
(b) $FP_{1}=0.1,FN_{1}=0.05,FP_{2}=0.5,FN_{2}=0.35$; Monopoly with larger $p_{1}^{*}=0.15$.
Figure 2: Price equilibria under strict domination with varying fixed FP and FN rate values.